# From the CDC to emerging infectious disease publics: The long-now of polarizing and complex health crises

### v2

@@ 5.1.3. What is in a mention?
    Pharmaceutical companies like @pfizer (specifically focused on the vaccine production)
@@
    Public health professionals/academics: @DrEricDing (see §2.1.3)
@@
    Politicians: @VP, @secazar, @joebiden, @govrondesantis, @mikepense, @presssec, @kamalaharris, @statedepartment, @nycgovcuomo,@joebiden, @gop, @statedept, @cdcdirector
  4. (4)
    News corporations and journalists: @foxnews, @jaketapper, @cnn, @abc, @cbsnews, @nbcnews @msnbc, @maddow, @alexberenson (this last account is a verified journalist who happens to share low-credibility information sources; more in §5.3)
  5. (5)
    State and federal agencies: @fema (Federal Emergency Management Agency), @nih, @us_fda (US Food and Drug Administration), @dhsgov (Department of Homeland Security), @hhsgov (health and human services), @cdcmmwr (Morbidity and Mortality Weekly Report), @drnancym_cdc, @HealthyFla (Florida Dept. of Health; see §6.2.1)
The results of our ERGM model, which analyzes the likelihood of a mention tie, are summarized in Table 7. Our edges (intercept) parameter is negative and significant. This means that there is a high ‘cost’ to the creation of a tie in the network (Keegan et al., 2012)[p.432]. In our sparse network, we found a positive “mutual”parameter suggesting a higher likelihood of tie creation associated with reciprocation. In other words, when one user mentions another, they are more likely to mention the original user.
Figure 4. Mention co-occurrence network showing top accounts mentioned along with the CDC account.
When analyzing node characteristics, we found that most of the discussion themes identified in §4.1.1 decreased the likelihood of creating a mention tie. The only theme that increased the likelihood of the creation of a mention tie was the one focused on non-COVID discourse probably due to the high connectivity between those working with opioid addiction patients, pain activists, and veterans (discussed in more detail in §6.3)). We found that out-degree tie formation was more likely for users who posted more using less credible sources. In other words, users who relied on less reputable sources were more likely to mention others while they themselves were less likely to be mentioned. Both negative and positive sentiment decreased the likelihood of the creation of mention ties, while neutral sentiment increased the likelihood. The more the use of rich media in Tweets, the less likely the creation of mention ties. Finally, verified users are less likely to have mention ties with other users.
Table 7. Exponential family random graph model of mentions network. A smaller value means the likelihood of creating a mention tie is smaller and vice-versa. **** p¡0.0001; *** p¡0.001; ** p¡0.01; * p¡0.05
Estimate Std. Error z value
nodeocov.verified -0.062 0.012 -4.979 ***
nodeicov.verified -0.205 0.013 -15.236 **
nodematch.top_theme.Mitigation -0.147 0.007 -20.031 ***
nodematch.top_theme.Non-COVID 1.236 0.032 38.792 ***
nodematch.top_theme.Politics -0.205 0.011 -18.738 ***
nodematch.top_theme.Test -0.267 0.022 -12.250 **
nodeocov.lowq-credibility 0.096 0.024 3.948 ***
nodeicov.lowq-credibility -0.231 0.029 -7.980 ***
nodeicov.photo -0.267 0.008 -32.107 ***
nodematch.sentiment.negative -0.237 0.009 -25.842 ***
nodematch.sentiment.neutral 0.381 0.011 33.583 ***
nodematch.sentiment.positive -0.078 0.008 -9.781 ***
edges -8.896 0.004 -2000.335 ***
mutual 3.483 0.063 55.150 ***
5.2. Sentiment, rich media, and low credibility sources
In this section, we will analyze the qualities of positive, neutral, and negative labeled Tweets, as well as ones labeled with the use of rich media (e.g., images) and low-credibility sources.
5.2.1. Positive Tweets
A higher retweet and quote Tweet count, being verified, and having a higher follower-following ratio are all more positively predictive of positive Tweets. Sharing data sources pointing to the seriousness of COVID (Topic 1), but also to good news about dropping rates of infection (Topic 15).
Positive Tweets are less likely to receive direct responses. They are also less likely to discuss COVID and seasonal flu infection and mortality rates (Topics 16 & 25), the efficacy of masks (Topics 2 and 54), or calling for obedience to the CDC mitigation guidelines in general (Topic 44). They tend to make less mention of political discourse like Trump’s CDC (Topic 8), the political/legislative response to COVID (Topic 29), or comparing the COVID response to H1N1 response under Obama’s administration (Topic 56).
5.2.2. Neutral Tweets
Higher follower-following ratio increases the chance of a neutral Tweet. Neutral tweets tend to receive fewer responses, quote-tweets, and likes.
Sports players and attendance guidelines (Topic 21) and social distancing guidelines (six feet apart; Topic 63) are also positively associated with neutral Tweets. These are often posts by users with high FFR, usually at a local level, like high school and college athletic departments, coaches, and sports commentators who, while not verified, have large followings. Many of the posts simply stated the current requirements for attending sports events (e.g., in-door masking for intramural sports or social distancing requirements).
The lack of clarity in CDC guidelines (Topic 9), the attack and defense of the CDC (Topic 14), and the administrative and legislative response to COVID (Topic 32) are positively predictive of neutral Tweets. While defending/attacking the CDC might sound polarizing, the Tweets themselves can have a more neutral sentiment since they can be nuanced. For example, the author of the Tweet could specify that while they support the CDC’s efforts, they do not think the organization is doing enough to counter the virus. Specifically, they could focus on the fact that the CDC only publishes data when it can have a more positive context – when the COVID infection numbers are decreasing. For reference, Table 1 shows that CDC Tweets are more positive than any other entity, cluster, or theme with an average VADER score of 0.17.
5.2.3. Negative Tweets
Higher retweets and follower-following ratios are less predictive of negative Tweets. The more a Tweet discusses COVID infections and mortality rates (Topic 25) or the opioid pandemic data (Topic 19), the more they are likely to evince a negative sentiment. The first topic can be more negative because it usually focuses on attacking the CDC for reporting mortality numbers that are unexpectedly high or low (note that while COVID infection numbers are heavily discussed in positive Tweets, positive Tweets do not focus on mortality rates). Negative tweets are also less focused on attacking/defending the CDC (Topic 14) and in-person gathering guidelines.
5.2.4. Tweets using rich media
Tweets using rich media are more likely to be retweeted, liked, quote-tweeted, get a reply, and engage with the CDC (Topic 5). Tweets using rich media are less likely to be posted from verified accounts, or to discuss lack of up-to-date COVID testing data (Topic 6), seasonal flu mortality rates (Topic 16) or music and stress (Topic 61).
5.2.5. Tweets using low credibility sources
The use of rich media and verified accounts are more predictive of the use of low credibility sources. The focus on discussing the opioid pandemic (Topic 19) and COVID mortality rates (Topic 7) are also positively predictive of low credibility sources.
Tweets using low credibility sources are less likely to be liked, retweeted, get responses, or engage in discussions regarding legislative and administrative responses to COVID (Topic 32) or Salmonella outbreaks (Topic 23).
5.3. Verified accounts and accounts with high follower-following ratios
Verified accounts are more likely to be retweeted, quote-tweeted, liked, get direct responses, use rich media and engage in attacking or defending the CDC (Topic 14). Verified accounts are less likely to Tweet in the CDC public (Tweet count) and less likely to discuss Sports players and attendance guidelines (Topic 21). While verified accounts engaged with less Tweets than other users, their messages were retweeted at a higher rate which is in-line with the high echo value for verified accounts indicated in Table 1. It is also interesting to note that responses to verified accounts in general, and the CDC in particular were around 5, which means that the user received 5 retweets for an engagement with a verified account or the CDC.
While the use of low-credibility sources is not predictive of verified accounts, verified accounts are predictive of the use of low credibility sources in a Tweet. In other words, some verified accounts used low credibility sources in their engagement. One example is @alexberenson, a journalist who pushed back against most of the CDC’s mitigation strategies. Berenson, with a verified journalist account, used several low-credibility sources and was mentioned along with the @CDCGov account with high frequency (see §5.1.3) which would expand his audience further.
Accounts with high FFR were more likely to receive likes and responses. They were more likely to engage in positive or neutral messaging and use rich media in their Tweets.
6. Study II: Discourse Themes, Interest Publics, and Stakeholder Groups
In this section, we answer RQ2 by analyzing framing used by different publics and how they are polarized, and we identify the main stakeholders engaging in each public.
Using the top 10 hashtags and the co-occurrence network of hashtags that are associated with #CDC (Türker and Sulak, 2018) (Figure 5), we found the following main areas of focus:
  1. (1)
    Mitigation:
    1. (a)
      Masking: #maskup, #masks, #wearamask
    2. (b)
      Vaccines: #getvaccinated, #vaccines,, #covidvaccine
    3. (c)
      Lock-downs: #quarantine, #socialdistancing, #stayhome
  2. (2)
    Science and medical expertise: #NIH, #WHO, #FDA, #DrFauci, Healthcare Common Procedure Coding System (#HCPCS): focusing on medical insurance
  3. (3)
    Political (Trump) and well-known public health figures (Dr. Fauci). #Florida is also the top-tenth hashtag due to the political fights between the Florida state government’s handling of the disease and the CDC guidelines (discussed in more detail in §6.2.1): #MAGA, #whitehouse, #trumpliesamericansdie, #trumpisaliar, #trumpvirus
Table 15 shows these top themes, along with their echo (total retweets/total number of tweets by user), the ratio of tweets by verified users, tweets with rich media, and low-credibility tweets. As shown in the table, all of the themes have a negative kurtosis, or wide distribution with more extreme scores than those in a normal distribution. This means that the discussions were polarized (Kallner, 2017). In this section, we analyze the mostly polarized discussions under each theme, and we introduce the different stakeholders and their respective positive, neutral, and negative communication frames.
Summary of findings:
We found that there are three main areas of discourse in the CDC public: (1) Mitigation guidelines: introducing the different CDC mitigation guidelines and how changes to them over time (not recommending masking for all at the beginning, then recommending it, then saying the vaccine will negate the need for masking, until the Omicron variant took hold); (2) COVID politics: how the CDC was not ready for the pandemic, and how the roll-out of the vaccine for the population was considered good progress after losing trust in the CDC early in the pandemic; and (3) Non-COVID discourse: while issues like Salmonella outbreaks and vaping dangers become important as COVID takes hold, some other conditions like opioid addiction continue to be public health issues now compounded by lock-downs and other COVID complications.
Table 8. Cluster top topics Part I
@@ 6. Study II: Discourse Themes, Interest Publics, and Stakeholder Groups
Table 9. Cluster top topics Part II
@@ 6.1.1. Mitigation guidelines in the social context
In Cluster 12, these concerns were localized to school district sports activities and how they can be reopened safely for families, teachers, and student-athletes. While some parents were pushing for a return to ”normal,” medical professionals pointed out that COVID infections of youth athletes can have long-term effects on them. They cautioned against reopening sporting events without adequate mitigation, including COVID testing. Many accounts included high school and college athletic departments (N=11), coaches and heads of athletic teams (N=5), and school administrators (N=2). A focus on relaying the latest mitigation guidelines by the CDC for sporting events was evident in neutral Tweets as discussed in §5.2.2.
@@ 6.1.2. As the pandemic evolves, guideline clarity devolves
Cluster 7 is focused on the idea that CDC guidelines are not clear. The discourse in this cluster was mostly pro-mitigation and leaned on the liberal side politically. For example, in Table 12, we can see that words like #ANTIFA, #BlackLivesMatter, etc. were indicative of the biographies in this cluster. The focus on professional status was also important in this cluster (see words like scholar, chief medical resident etc.). Medical professionals (N=17) wanted the CDC to issue simpler guidelines and not change them in response to subtle changes in the number of COVID infections. Many also pointed out that the CDC should use other metrics besides death rates because the medical and scientific community still has little understanding of long-term effects of COVID infections. The change in quarantine guidelines was critiqued by vascular and pediatric medical doctors in Cluster 16. Specifically, they were concerned about the shortening of the quarantine period from 14 days to five days. This issue was also picked up by different stakeholders in Cluster 13. While reporters were presenting the CDC guideline changes, epidemiologists (N=4) and other healthcare professionals (N=8) as well as medical institutions (N=5) argued against this change. They believed five days did not provide enough protection against spreading COVID to others. However, this was countered by other stakeholders, including the Air Line Pilots Association and travelers, who professed to have lost patience with all the new travel guidelines. In a related way, many posters were excited about CDC guidelines that do not require tests or masking on airplanes or indoors if fully vaccinated. Some medical professionals (N=5) and advocates (N=4) still encouraged masking to protect themselves and others. There was disappointment in the new guidelines and worry among essential workers about un-vaccinated individuals potentially exposing them to the virus. As masking guidelines were changing, questions arose about the efficacy of the vaccine in stopping infections. These questions emerged especially as the vaccine became less effective after the Delta and Omicron variants (Cluster 4). Some of the discourse touched on reintroducing masking, especially at peak COVID infection points (e.g., the Omicron wave in late 2021). Another point of friction focused on the changes that came with re-opening schools. Both are discussed in more detail in §7.1.
@@ 6.1.3. School management post-COVID
Cluster 15 is focused on school openings. The discussions in this cluster are split between those pushing for opening schools back up (N=10), three of whom self-identified as parent advocates. One of the parent advocates argue school closures have negatively affected the managing of adverse childhood experiences (ACE) since many of the mandatory reporters (e.g., social workers, teachers, school councilors, etc.) do not have contact with children. Teachers (N=7), superintendents (N=3), and teacher union leaders (N=1) called for teachers to have a higher priority on the vaccination list to potentially allow for faster school re-openings, which made many parents happy. However, since the vaccines were not approved for children at the time, those same accounts were worried about COVID being transmitted between students. Therefore, these accounts called for masking in schools regardless of vaccination status. At the same time, many of the advocate parents that were introduced earlier also perceived masking as unsafe and unfair to children. Indeed, we found that school re-openings and mitigation guidelines can be polarizing and engaging more retweets (§5.1.1) and responses (§5.1.2).
@@ 6.2.1. Test-trace and mortality rates
The lack of COVID test kits, especially in the early days of the pandemic, is central to this theme. As seen in Table 12, the accounts of public health professionals (N=5), medical research institutions (N=2), and self-identified liberal politicians (N=3) engaged in critiquing, and at times directly attacking, the Trump administration’s management of testing. The main issue they focused on was that accurate reflection of infection counts is important for safeguarding others, especially in healthcare settings (e.g., #nurselife). They used hashtags like #wherearethetests, #liesaredeadly, and #incompetentinchief. The focus on the lack of testing and questioning mortality rates was associated with negative Tweets as discussed in §5.2.3. There was a specific focus on the COVID testing strategy in Florida where the @HealthyFla (Florida Dept. of Health) determined that only symptomatic people would be tested for COVID as opposed to testing a larger sample of the population to assess whether lock-downs are needed. Governor Desantis’ account (@govrondesantis) is frequently mentioned in this discussion as many point to his influence in this COVID testing strategy. An example of rich media messages about the topic can be seen in Figure 22. Given how politicized the testing theme has been, we will explore politicization of the discourse further next.

### v3

@@ 5. Findings
    Mitigation – including hashtags related to masking (#maskup, #masks), vaccination (#getvaccinated, #covidvaccine), and lockdowns (#quarantine, #stayhome).
@@
    Science and medical expertise – featuring #NIH, #WHO, #FDA, #DrFauci, and #HCPCS (Healthcare Common Procedure Coding System), the latter linked to insurance-related discourse.
@@
    Politics and public figures – centered on Trump and Fauci, with #Florida prominent due to state-versus-CDC conflicts (#MAGA, #trumpvirus, etc.; see more detail in §5.4.1).
We now explore these discourse strands across different publics in more detail next.
5.1. Many echo-chambers
The E-I index analysis reveals distinct patterns of interaction across the 16 discourse clusters. Clusters 5 through 16 each have an E-I index of -1.0, indicating perfect echo chambers—users in these clusters interacted exclusively with others who shared the same stance, with no cross-stance engagement. Cluster 12 stands out with a moderately negative E-I index of -0.24, reflecting some interaction across stance lines but still largely internal communication. Similarly, Cluster 3 shows a slight echo chamber tendency with an E-I index of -0.19. In contrast, Clusters 2 and 1 exhibit the most cross-cutting discourse: Cluster 2 has a strongly positive E-I index of 0.75, while Cluster 1 is the most integrative, with an E-I index of 1.0, meaning all interactions occurred between users of different stances. Clusters 4, 10, and 14 could not be evaluated due to missing or insufficient interaction data. These results suggest that while many clusters are ideologically siloed, a few serve as bridging spaces where users engage across lines of disagreement.
Table 1. Top topics and theme distributions for each discourse cluster, including major focus areas and retweet averages (part I)
@@ 5.1. Many echo-chambers
Table 2. Top topics and theme distributions for each discourse cluster, including major focus areas and retweet averages (part II)
@@ 5.2.1. Mitigation guidelines in the social context
In Cluster 12, these concerns were localized to school district sports activities and how they can be reopened safely for families, teachers, and student-athletes. While some parents were pushing for a return to ”normal,” medical professionals pointed out that COVID infections of youth athletes can have long-term effects on them. They cautioned against reopening sporting events without adequate mitigation, including COVID testing. Many accounts included high school and college athletic departments (N=11), coaches and heads of athletic teams (N=5), and school administrators (N=2).
@@ 5.2.2. As the pandemic evolves, guideline clarity devolves
Cluster 7 is focused on the idea that CDC guidelines are not clear. The discourse in this cluster was mostly pro-mitigation and leaned on the liberal side politically. For example, in Table 3, we can see that words like #ANTIFA, #BlackLivesMatter, etc. were indicative of the biographies in this cluster. The focus on professional status was also important in this cluster (see words like scholar, chief medical resident etc.). Medical professionals (N=17) wanted the CDC to issue simpler guidelines and not change them in response to subtle changes in the number of COVID infections. Many also pointed out that the CDC should use other metrics besides death rates because the medical and scientific community still has little understanding of long-term effects of COVID infections. The change in quarantine guidelines was critiqued by vascular and pediatric medical doctors in Cluster 16. Specifically, they were concerned about the shortening of the quarantine period from 14 days to five days. This issue was also picked up by different stakeholders in Cluster 13. While reporters were presenting the CDC guideline changes, epidemiologists (N=4) and other healthcare professionals (N=8) as well as medical institutions (N=5) argued against this change. They believed five days did not provide enough protection against spreading COVID to others. However, this was countered by other stakeholders, including the Air Line Pilots Association and travelers, who professed to have lost patience with all the new travel guidelines. In a related way, many posters were excited about CDC guidelines that do not require tests or masking on airplanes or indoors if fully vaccinated. Some medical professionals (N=5) and advocates (N=4) still encouraged masking to protect themselves and others. There was disappointment in the new guidelines and worry among essential workers about un-vaccinated individuals potentially exposing them to the virus. As masking guidelines were changing, questions arose about the efficacy of the vaccine in stopping infections. These questions emerged especially as the vaccine became less effective after the Delta and Omicron variants (Cluster 4). Some of the discourse touched on reintroducing masking, especially at peak COVID infection points (e.g., the Omicron wave in late 2021). Another point of friction focused on the changes that came with re-opening schools.
@@ 5.2.3. School management post-COVID
Cluster 15 is focused on school openings. The discussions in this cluster are split between those pushing for opening schools back up (N=10), three of whom self-identified as parent advocates. One of the parent advocates argue school closures have negatively affected the managing of adverse childhood experiences (ACE) since many of the mandatory reporters (e.g., social workers, teachers, school councilors, etc.) do not have contact with children. Teachers (N=7), superintendents (N=3), and teacher union leaders (N=1) called for teachers to have a higher priority on the vaccination list to potentially allow for faster school re-openings, which made many parents happy. However, since the vaccines were not approved for children at the time, those same accounts were worried about COVID being transmitted between students. Therefore, these accounts called for masking in schools regardless of vaccination status. At the same time, many of the advocate parents that were introduced earlier also perceived masking as unsafe and unfair to children.
@@
5.3. Non-COVID CDC guidelines
Table 3. Cluster metadata including user counts, verified status, top biography descriptors (doc2vec and LLR), and E-I index. These indicators reflect ideological cohesion and engagement diversity within each cluster (part I).
# Cluster Name Users Top desc doc2vec Top desc LLR E-I Index
1 COVID racial disparities 121 (13) @georgetown, displaced, self-proclaimed, care, oncology, #healthequity, alliance, wellness, Latinx, #Harmreduction, chicano, #Goodtrouble, ADOS law, latino, black, program, medicine, POTUS, community, gender, socially, racist, violence, getvaccinated, epidemiologist, care, feminist, adjunct, diversity, researcher 1.00
2 Mitigation policy discussion 39,749 (1,932) first amendment,#ANTIFA, #BuildTheWall, #ClimateAction, #CONSERVATIVE, #RightMatters, #AllLivesMatter, #Metoo, #MAGA, Supporter, #Truthmatters freedoms, dweller, proponent, bioethics, thedemocrats, Johns Hopkins University, Come get them 0.75
3 Partisan discussion of COVID response 4,918 (93) #patriot, #bidenharris2020,#independent, GoBucks, #StrongerTogether, #Science, Realist, #TeamPelosi, Latter-Day, #BlueWave, #VoteBlueNoMatterWho retired, trump, resist, donald, steelernation, detest, theresistance, equalrights, hamiltonmusical, election, angry -0.19
4 Guideline changes with Omicron Variant 41(6) Educator, @Harvard, Educated, #books, Forbes, @ClemsonUni, lecturer, mgmt, Epidemiology, Strategist, CEO, @HowardU njspotlightnews, action10news, cichealth, byline, columbiamsph, kennedy_school, columbiamed, newscaster, nyuniversity
5 CDC 1(1) NA privacy, centers, cdc, updates, source, control, prevention, daily, disease, safety, official, health -1.00
6 Pro-mitigation 91(1) #Geek, OPINIONS, Makeup, Thinker, #DubNation, #RedSox, #Braves long distance relationship, nonconformist, hospitalityindustry, card -1.00
7 No clarity in CDC guidelines 1,280(40) scholar, #ANTIFA, hacker, #SaveOurDemocracy, #BlackTransLivesMatter fanboys, relax, wtvc, dogfather, chiefmedicalresident, finanacialprofessional -1.00
@@ 5.4.1. Test-trace and mortality rates
The lack of COVID test kits, especially in the early days of the pandemic, is central to this theme. As seen in Table 3, the accounts of public health professionals (N=5), medical research institutions (N=2), and self-identified liberal politicians (N=3) engaged in critiquing, and at times directly attacking, the Trump administration’s management of testing. The main issue they focused on was that accurate reflection of infection counts is important for safeguarding others, especially in healthcare settings (e.g., #nurselife). They used hashtags like #wherearethetests, #liesaredeadly, and #incompetentinchief. The focus on the lack of testing and questioning mortality rates was associated with negative Tweets as discussed in §LABEL:subsec:negative. There was a specific focus on the COVID testing strategy in Florida where the @HealthyFla (Florida Dept. of Health) determined that only symptomatic people would be tested for COVID as opposed to testing a larger sample of the population to assess whether lock-downs are needed. Governor Desantis’ account (@govrondesantis) is frequently mentioned in this discussion as many point to his influence in this COVID testing strategy. Given how politicized the testing theme has been, we will explore politicization of the discourse further next.